OpenStudy (anonymous):

If a = 4 and b = 5, what is the value of the following expression? 24 ÷ a + 2 • b 8 40 14 16

OpenStudy (anonymous):

@DJBreezy

OpenStudy (djbreezy):

The answer is 16

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

thank u so much

OpenStudy (anonymous):

24÷4+2*5 =6+2*5 =6+10 =16

OpenStudy (anonymous):

thank u @ImJRF

OpenStudy (anonymous):

your welcome

OpenStudy (djbreezy):

medal me and i'll medal @ImJRF